OpenStudy (anonymous):

X+Y+2=36 4X+2Y+Z=71 6X+5Y+4Z=85 can someone help me solve this equation? I NEED help. Anyone

OpenStudy (istim):

What do you mean by 'solve'?

OpenStudy (istim):

And please don't repost.

OpenStudy (anonymous):

If someone can help me step by step. I really do need help.

OpenStudy (istim):

Solve for what, x or y?

OpenStudy (istim):

Wait, I think Bookworm got this.

OpenStudy (anonymous):

Is the 2 in the first equation supposed to be a z?

OpenStudy (anonymous):

Yes, I had just noticed that.

OpenStudy (anonymous):

I had got 93 for X but I'm not sure if it's correct..

OpenStudy (anonymous):

Alright. So start by multiplying the top equation by -4. After you do that, set that new equation above the second equation. The -4x and 4x cancel out. Add the remaining values (y and the constants) Then divide the coefficient of y by the sum of the constants. Repeat so on to solve for x,y, and z.

OpenStudy (anonymous):

Thank you, I'll try solving this but I'm not 100% sure about getting it right.

OpenStudy (anonymous):

so, after I cancel the 4's will it look like -2y+9=107? @AlwaystheBookworm

OpenStudy (anonymous):

hey, dont be doubtful just find ur answers and put them in any equation and varify that if both sides of equation are equal ur answers will be right then...........

OpenStudy (anonymous):

Ahhh, It's cause this is so confusing! :/ @khurramshahzad

OpenStudy (anonymous):

You should be getting -2y-3z=-73 Here's how it works: -4(x+y+z=36) -4x-4y-4x=-144 Set that to 4x+2y+z=71 So you get: -4x-4y-4z=-144 4x+2y+z=71 Now cancel out the x and add the y,z, and constant. You should get: -2y-3z=-73

OpenStudy (anonymous):

u can varify it ur all confusion will vanish if u go and varify . it will be clear that weather ur question is true of false k

OpenStudy (anonymous):

Beware, the answers are fractions, not integers.

OpenStudy (anonymous):

Ohh, I see. Then what do I do afterwards? @AlwaystheBookworm @robtobey I didn't know that can be possible..

OpenStudy (anonymous):

Take your first equation again. Instead of setting it to the second equation, set it to the third one. x+y+z=36 6x+5y+4z=85 Now, eliminate the x just like in the previous system. The only difference is you would multiply the top by -6, instead of -4.

OpenStudy (anonymous):

Trust, but verify.\[\left\{X=\frac{97}{7},Y=\frac{141}{7},Z=-\frac{173}{7}\right\} \]

OpenStudy (anonymous):

gg